# Kolmogorov-Arnold Networks (KANs) for Time Series Analysis

### v1

@@ Abstract
Index Terms—  Kolmogorov-Arnold Networks, ML, Time series
1 Introduction
Time series forecasting plays a key role in a wide range of fields, driving critical decision-making processes in finance, economics, medicine, meteorology, and biology, among others, reflecting the wide applicability and its significance across many domains. It involves predicting future values based on the previously observed data points. With this goal in mind, understanding the dynamics of time-dependent phenomena is essential and requires unveiling the patterns, trends and dependencies hidden with the historical data. While conventional approaches have been traditionally centered on parametric models grounded in domain-specific knowledge, such as autoregressive (AR), exponential smoothing, or structural time series models, contemporary Machine Learning (ML) techniques offered a pathway to discern temporal patterns solely from data-driven insights.
@@ 1 Introduction
More recently, ML models have significantly impacted the forecasting landscape by handling large datasets and capturing complex nonlinear relationships that traditional methods cannot. In recent years, Deep Learning (DL)-based forecasting models have gained popularity, motivated by the notable achievements in many fields. For instance, neural networks have been extensively studied due to their flexibility and adaptability. Simple Multi-Layer Perceptron (MLPs) were among the first to be applied to forecasting problems, demonstrating significant potential in non-linear data modeling [5].
@@
Built upon these light models, more complex architectures have progressively expanded the capabilities of neural networks in time series forecasting. Typical examples are recurrent neural network architectures such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs), which are designed to maintain information in memory for long periods without the risk of vanishing gradients – a common issue in traditional recurrent networks [6]. On a related note, Convolutional Neural Networks (CNNs), which are fundamentally inspired by MLPs, are also extensively employed in time series forecasting. These architectures are particularly efficient at processing temporal sequences due to their strong spatial pattern recognition capabilities. The combination of CNNs with LSTMs has resulted in models that efficiently process both spatial and temporal dependencies, enhancing forecasting accuracy [7]. These models have started to outperform established benchmarks in complex forecasting tasks, motivating a significant shift towards more complex network structures. Unfortunately, as the majority of the models mentioned above are inspired by MLP architecture, they tend to have poor scaling law, i.e., the number of parameters in MLPs networks do not scale linear with the number of layers, and often lack interpretability.
@@
A recent study in reference [8], which was presented just 12 days prior to the submission of this paper, introduces Kolmogorov-Arnold Networks (KANs), a novel neural network architecture designed to potentially replace traditional multilayer perceptrons. KANs represent a disruptive paradigm shift, and as a potential game changer have recently attracted the interest of the AI community worldwide. They are inspired by the Kolmogorov-Arnold representation theorem [9], [10]. Unlike MLPs, which are inspired by the universal approximation theorem, KANs take advantage of this representation theorem to generate a different architecture. They innovate by replacing linear weights with spline-based univariate functions along the edges of the network, which are structured as learnable activation functions. This design not only enhances the accuracy and interpretability of the networks, but also enables them to achieve comparable or superior results with smaller network sizes across various tasks, such as data fitting and solving partial differential equations. While KANs show promise in improving the efficiency and interpretability of neural network architectures, the study acknowledges the necessity for further research into their robustness when applied to diverse datasets and their compatibility with other deep learning architectures. These areas are crucial for understanding the full potential and limitations of KANs.
@@
Our paper is a prospective study that investigates the application of KANs to time series forecasting. To the best of authors’ knowledge, not previously explored in the literature. We aim to evaluate the practicality of KANs in real-world scenarios, analyzing their efficiency in terms of the number of trainable parameters and discussing how the additional degrees of freedom might affect forecasting performance. Herein, we will assess the performance using real-world satellite traffic data. This exploration seeks to further validate KANs as a versatile tool in advanced neural network design for time series forecasting, although more comprehensive studies are required to optimize their use across broader applications. Finally, we note that due to the early stage of KANs, it is fair to compare it as a potential alternative to MLPs, but further investigation is needed to develop more complex solutions that can compete with advanced architectures such as LSTMs, GRUs and CNNs.
@@
This paper is structured as follows. Section 2 presents the problem statement, providing fundamental background on Kolmogorov-Arnold representation theorem and KANs. Section 3 will present the experimental setup description. Simulation results analyzing the performance of KANs with real-world datasets are shown in Section 4. Finally, concluding remarks are provided in Section 5.
2 Problem statement
@@
2.1 Kolmogorov-Arnold representation background
@@
2.2 Kolmogorov-Arnold network background
@@
Table 1: Model configurations for satellite traffic forecasting
@@
A KAN layer is defined by a matrix $\mathbf{\Phi}$ [8] composed by univariate functions $\{\phi_{i,j}(\cdot)\}$ with $i=1,...,N_{in}$ and $j=1,...,N_{out}$, where $N_{in}$ and $N_{out}$ denote the number of inputs and the number of outputs, respectively, and $\phi_{i,j}$ are the trainable spline functions described above. Note according to the previous definition, the Kolmogorov-Arnold representation theorem presented in Section 2.1 can be expressed as a two-layer KAN. The inner functions constitute a KAN layer with $N_{in}=n$ and $N_{out}=2n+1$, while the external functions constitute another KAN layer with $N_{in}=2n+1$ and $N_{out}=1$.
@@
Fig. 1: Example of normalized satellite traffic series data with the conditioning and prediction lengths denoted in blue, and red, respectively.
2.3 KAN time series forecasting network
@@
$\mathbf{y}=\text{KAN}(\mathbf{x})=(\mathbf{\Phi}_{2}\circ\mathbf{\Phi}_{1})\mathbf{x}.$ (5)
where the output functions $\Phi_{2}$ generates the $N_{o}$ outputs values corresponding to (1) by doing the transformation from the previous layers. The proposed network can be used to forecast future traffic data in the prediction length, based solely on the context length. Fig. 2 shows a generic representation for any arbitrary number of layers $L$.
Fig. 2: Example of the flow of information in the KAN network architecture for our traffic forecasting task. Learnable activations are represented inside a square box.
3 Experimental setup
The dataset has been generated within the context of the European project 5G-STARDUST. The inputs are obtained from a satellite operator (SO), as a result of processing real information from a GEO satellite communication system, which provisions broadband services. The dataset is a long time series capturing aggregated traffic data. To preserve privacy, anonymous clients have been defined with more than 500 connected users, and the traffic has been normalized. The measurements are monthly long, and the time granularity is 1 hour.
The traffic has been extracted per satellite beam in Megabits per second (Mbps). Although the data has been collected using a GEO satellite communication system, it is expected that user needs could be used to address LEO systems, as well. It is worth emphasizing that the data collected can be used for AI-driven predictive analysis, to forecast traffic conditions, which is essential to avoid congestion and to make efficient use of satellite resources. Endowing the network with intelligence will be beneficial to meet the different demands of satellite applications.
@@ 3 Experimental setup
Fig. 3: Satellite traffic over three different beams with their forecasted values using a 4-depth KAN and a 4-depth MLP.
4 Simulation results
This section investigates the forecasting performance of different KAN and MLP architectures for predicting satellite traffic over the six beam areas. Concretely, we have a context length of 168 hours (one week) and a prediction length of 24 hours (one day). This translates to $T=24$, $c=168$, where $y_{t_{0}+T}=192$ in (1) and (2). Our focus is on evaluating the efficacy of KAN models compared to traditional MLPs. We designed our experiments to compare models with similar depths but varying architectures to analyze their impact on forecasting accuracy and parameter efficiency. Table 1 summarizes the parameters selected for this evaluation. We have data for the six beams over one month. We use two weeks + 1 day for training and one week + 1 day for testing for all beams that were not seen by the network. We train all the networks with $500$ epochs and Adam optimizer with a learning rate of $0.001$. The selected loss function minimizes the mean absolute error (MAE) of the values around the prediction length.
4.1 Performance analysis
We analyze the forecasting performance in the prediction length. Figures 3a-c depicts the real traffic value used as input (in green) to the networks, the expected output prediction length (in blue) and the values predicted values using a KAN (in red) and MLP (in purple) of depth 4 both – see Table 1 for details on model configuration. In general, our results show that the predictions obtained using KANs better approximates the real traffic values than the predictions obtained using traditional MLPs.
@@ 4.1 Performance analysis
Furthermore, the parameter count reveals a significant difference in model complexity. KAN models are notably more parameter-efficient, with KAN (4-depth) utilizing only 109k parameters compared to 329k parameters for MLP (4-depth) or 238k for MLP (3-depth). This reduced complexity suggests that KANs can achieve higher or comparable forecasting accuracy with simpler and potentially faster models. Such efficiency is especially valuable in scenarios where computational resources are limited or where rapid model deployment is required. The results also show that with an augmentation of 16k parameters in KAN, the performance can be significantly improved, contrary to MLPs which an increment of 91k parameters does not showcase a significant improvement.
@@
Table 2: Results summary
@@
4.2 KANs parameter-specific analysis
@@
Fig. 4: Ablation comparison of KAN-specific parameters.

### v2

@@ Abstract
Index Terms:
Kolmogorov-Arnold Networks, ML, Time series, Satellite
I Introduction
Time series forecasting is a traditional problem that plays a key role in a wide range of fields, driving critical decision-making processes in finance, economics, medicine, meteorology, and biology, reflecting the wide applicability and its significance across many domains [1, 2, 3, 4]. It involves predicting future values based on the previously observed data points. With this goal in mind, understanding the dynamics of time-dependent phenomena is essential and requires unveiling the patterns, trends and dependencies hidden with the historical data. While conventional approaches have been traditionally centered on parametric models grounded in domain-specific knowledge, such as autoregressive (AR), exponential smoothing, or structural time series models, contemporary Machine Learning (ML) techniques offered a pathway to discern temporal patterns solely from data-driven insights.
@@ I Introduction
More recently, ML models have significantly impacted the forecasting landscape by handling large datasets and capturing complex nonlinear relationships that traditional methods cannot. In recent years, Deep Learning (DL)-based forecasting models [9, 10] have gained popularity, motivated by the notable achievements in many fields. For instance, neural networks have been extensively studied due to their flexibility and adaptability. Simple Multi-Layer Perceptron (MLPs) were among the first to be applied to forecasting problems, demonstrating significant potential in non-linear data modeling [11, 3].
@@
Built upon these light models, more complex architectures have progressively expanded the capabilities of neural networks in time series forecasting. Typical examples are recurrent neural network architectures such as Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs), which are designed to maintain information in memory for long periods without the risk of vanishing gradients – a common issue in traditional recurrent networks [12, 13]. On a related note, Convolutional Neural Networks (CNNs), which are fundamentally inspired by MLPs, are also extensively employed in time series forecasting. These architectures are particularly efficient at processing temporal sequences due to their strong spatial pattern recognition capabilities. The combination of CNNs with LSTMs has resulted in models that efficiently process both spatial and temporal dependencies, enhancing forecasting accuracy [14]. These models have started to outperform established benchmarks in complex forecasting tasks, motivating a significant shift towards more complex network structures. Unfortunately, as the majority of the models mentioned above are inspired by MLP architecture, they tend to have poor scaling law [15], i.e., the number of parameters in MLPs networks do not scale linear with the number of layers, and often lack interpretability.
@@
A recent study in reference [16], which caught the attention of the research community, introduces Kolmogorov-Arnold Networks (KANs), a novel neural network architecture designed to potentially replace traditional multilayer perceptrons. KANs represent a disruptive paradigm shift, and as a potential game changer have recently attracted the interest of the AI community worldwide. They are inspired by the Kolmogorov-Arnold representation theorem [17, 18, 19]. Unlike MLPs, which are inspired by the universal approximation theorem, KANs take advantage of this representation theorem to generate a different architecture. They innovate by replacing linear weights with spline-based univariate functions along the edges of the network, which are structured as learnable activation functions. This design not only enhances the accuracy and interpretability of the networks, but also enables them to achieve comparable or superior results with smaller network sizes across various tasks, such as data fitting and solving partial differential equations. While KANs show promise in improving the efficiency and interpretability of neural network architectures, the study acknowledges the necessity for further research into their robustness when applied to diverse datasets and their compatibility with other deep learning architectures. These areas are crucial for understanding the full potential and limitations of KANs.
@@
Our paper is a prospective study that investigates the application of KANs to time series forecasting. We aim to evaluate the practicality of KANs in real-world scenarios, to the best of the authors’ knowledge, not previously explored in the literature, analyzing their efficiency regarding the number of trainable parameters and discussing how the additional degrees of freedom might affect forecasting performance. Herein, we will assess the performance using real-world satellite traffic data. This exploration seeks to further validate KANs as a versatile tool in advanced neural network design for time series forecasting, although more comprehensive studies are required to optimize their use across broader applications. Finally, we note that due to the early stage of KANs, it is fair to compare them as a potential alternative to MLPs, but further investigation is needed to develop more complex solutions that can compete with advanced architectures such as LSTMs, GRUs, and CNNs, already well-established on the MLP-based architectures [20, 21].
@@
This paper is structured as follows. Section 2 presents the problem statement, providing fundamental background on the Kolmogorov-Arnold representation theorem and describes our generalized KANs for time series forecasting. Section 3 introduces the experimental setup description. Simulation results analyzing the performance of KANs with real-world datasets are shown in Section 4. Finally, concluding remarks are provided in Section 5.
II Problem statement
@@
II-A Kolmogorov-Arnold representation background
@@
II-B Kolmogorov-Arnold network background
@@
TABLE I: Model configurations for satellite traffic forecasting
@@
A KAN layer is defined by a matrix $\mathbf{\Phi}$ [16] composed by univariate functions $\{\phi_{i,j}(\cdot)\}$ with $i=1,...,N_{in}$ and $j=1,...,N_{out}$, where $N_{in}$ and $N_{out}$ denote the number of inputs and the number of outputs, respectively, and $\phi_{i,j}$ are the trainable spline functions described above. Note according to the previous definition, the Kolmogorov-Arnold representation theorem presented in Section II-A can be expressed as a two-layer KAN. The inner functions constitute a KAN layer with $N_{in}=n$ and $N_{out}=2n+1$, while the external functions constitute another KAN layer with $N_{in}=2n+1$ and $N_{out}=1$.
@@
Figure 1: Example of normalized satellite traffic series data with the conditioning and prediction lengths denoted in blue, and red, respectively.
II-C KAN time series forecasting network
@@
$\mathbf{y}=\text{KAN}(\mathbf{x})=(\mathbf{\Phi}_{2}\circ\mathbf{\Phi}_{1})\mathbf{x},$ (5)
where the output functions $\Phi_{2}$ generates the $N_{o}$ outputs values corresponding to (1) by doing the transformation from the previous layers, i.e, we predict the $T$ time steps. The proposed network can be used to forecast future traffic data in the prediction length, based solely on the context length.
Fig. 2 shows a generic representation for any arbitrary number of layers $L$ as presented in (4).
Figure 2: Example of the flow of information in the KAN network architecture for our traffic forecasting task. Learnable activations are represented inside a square box.
III Experimental setup
The dataset has been generated within the context of the European project 5G-STARDUST. The inputs are obtained from a satellite operator (SO), as a result of processing real information from a GEO satellite communication system, which provisions broadband services. The dataset is a long time series capturing aggregated traffic data. To preserve privacy, anonymous clients have been defined with more than 500 connected users, and the traffic has been normalized. The measurements are monthly long, and the time granularity is 1 hour. The traffic has been extracted per satellite beam in Megabits per second (Mbps). Although the data has been collected using a GEO satellite communication system, it is expected that user needs could be used to address LEO systems, as well. It is worth emphasizing that the data collected can be used for AI-driven predictive analysis, to forecast traffic conditions, which is essential to avoid congestion and to make efficient use of satellite resources. Endowing the network with intelligence will be beneficial to meet the different demands of satellite applications.
We aim to investigate the forecasting performance of different KAN and MLP architectures for predicting satellite traffic over a total of six beam areas. Concretely, we have a context length of 168 hours (one week) and a prediction length of 24 hours (one day). This translates to $T=24$, $c=168$, where $y_{t_{0}+T}=192$ in (1) and (2). Our focus is on evaluating the efficacy of KAN models compared to traditional MLPs [^1] As KANs are in their infancy, we remark this comparison is fair instead of comparing against more complex architectures as LSTM.. We designed our experiments to compare models with similar depths but varying architectures to analyze their impact on forecasting accuracy and parameter efficiency. Table I summarizes the parameters selected for this evaluation. We have data for the six beams over one month. We use two weeks + 1 day for training and one week + 1 day for testing for all the different beams on the dataset. These test series were not seen by the network during training time. We train all the networks with $500$ epochs and Adam optimizer with a learning rate of $0.001$. The selected loss function minimizes the mean absolute error (MAE) of the values around the prediction length.
@@ III Experimental setup
Figure 3: Satellite traffic over three different beams with their forecasted values using a 4-depth KAN and a 4-depth MLP.
IV Simulation results
IV-A Performance analysis
We analyze the forecasting performance in the prediction length for different beams over the test set. Figures 3a-c depicts the real traffic value used as input (in green) to the networks, the expected output prediction length (in blue) and the values predicted values using a KAN (in red) and MLP (in purple) of depth 4 both – see Table I for details on model configuration. In general, our results show that the predictions obtained using KANs better approximates the real traffic values than the predictions obtained using traditional MLPs.
@@ IV-A Performance analysis
Furthermore, the parameter count reveals a significant difference in model complexity. KAN models are notably more parameter-efficient, with KAN (4-depth) utilizing only 109k parameters compared to 329k parameters for MLP (4-depth) or 238k for MLP (3-depth). This reduced complexity suggests that KANs can achieve higher or comparable forecasting accuracy with simpler and potentially lighter models. Such efficiency is especially valuable in scenarios where computational resources are limited or where rapid model deployment is required. The results also show that with an augmentation of 16k parameters in KAN, the performance can be significantly improved, contrary to MLPs which an increment of 91k parameters does not showcase a significant improvement.
@@
TABLE II: Results summary
@@
IV-B KANs parameter-specific analysis
@@
Figure 4: Ablation comparison of KAN-specific parameters
during training time.